Haas Company manufactures and sells one product. The following information pertains to each of the company’s first three years of operations: Variable costs per unit: Manufacturing: Direct materials $ 21 Direct labor $ 13 Variable manufacturing overhead $ 8 Variable selling and administrative $ 1 Fixed costs per year: Fixed manufacturing overhead $ 600,000 Fixed selling and administrative expenses $ 240,000 During its first year of operations, Haas produced 60,000 units and sold 60,000 units. During its second year of operations, it produced 75,000 units and sold 50,000 units. In its third year, Haas produced 40,000 units and sold 65,000 units. The selling price of the company’s product is $57 per unit.

Factory electricity costs for the last 3 years have been as follows:

Year 1:                   $40,000

Year 2:                   $45,000

Year 3:                   $28,000

a) Use the 3 years of electricity expense given above and YOUR units produced. Tells you the number of units produced/sold for the first, second and third years of operations. You should use this data with the High-Low method to calculate the variable cost per unit for electricity. b) Continue to use the High-Low method, identify the average fixed cost element for electricity. c) Write the formula for electric costs as Y = Fixed costs + VC*X. Your formula should include Y and X as variables and your answer to c) as Fixed Costs and your answer to d) as the VC per unit. e) If Haas Company plans to produce and sell 150,000 units, how much can they expect their electric costs to be for the period?
Answer #1

a.

Variable cost per unit for electricity = $45,000-28,000 / 75,000-40,000

Variable cost per unit for electricity = $17,000 / 35,000 = $0.486 per unit of electricity

b.

Fixed cost = $45,000 - 36,450 (75,000*$0.486) = $8,550

c.

Y = $8,550 + $0.486*X

e.

Y = $8,550 + $0.486*150,000

Y = $8,550 + 72,900 = $81,450

Expected electric costs for the period = $81,450